# NurValues: Real-World Nursing Values Evaluation for Large Language Models in Clinical Context

### v1

@@ Abstract
This work introduces the first benchmark for nursing value alignment, consisting of five core value dimensions distilled from international nursing codes: Altruism, Human Dignity, Integrity, Justice, and Professionalism. The benchmark comprises 1,100 real-world nursing behavior instances collected through a five-month longitudinal field study across three hospitals of varying tiers. These instances are annotated by five clinical nurses and then augmented with LLM-generated counterfactuals with reversed ethic polarity. Each original case is paired with a value-aligned and a value-violating version, resulting in 2,200 labeled instances that constitute the Easy-Level dataset. To increase adversarial complexity, each instance is further transformed into a dialogue-based format that embeds contextual cues and subtle misleading signals, yielding a Hard-Level dataset. We evaluate 23 state-of-the-art (SoTA) LLMs on their alignment with nursing values. Our findings reveal three key insights: (1) DeepSeek-V3 achieves the highest performance on the Easy-Level dataset (94.55), where Claude 3.5 Sonnet outperforms other models on the Hard-Level dataset (89.43), significantly surpassing the medical LLMs; (2) Justice is consistently the most difficult nursing value dimension to evaluate; and (3) in-context learning significantly improves alignment. This work aims to provide a foundation for value-sensitive LLMs development in clinical settings [^1] The dataset and the code are available at https://huggingface.co/datasets/Ben012345/NurValues.
@@ 1 Introduction
Like a powerful tide, large language models (LLMs) have swept across the landscape of natural language processing (NLP), demonstrating remarkable performance across a wide range of downstream tasks [49]. This transformative wave has rapidly propagated into vertical domains, notably healthcare [39]. Through continued pre-training on extensive medical knowledge repositories, a constellation of medical LLMs has emerged, such as Med-PaLM 2 [37], OpenBioLLM [6], LLaVA-Med [26], HuatuoGPT [11]. Such models demonstrate impressive capabilities in patient-facing Q&A, clinical reasoning, and even medical licensing examinations.
However, these advances have also surfaced critical risks, including hallucinated diagnoses, culturally insensitive or biased recommendations, and failures to respect patient autonomy. Given the high stakes of medical decision-making and its direct impact on patient well-being, there is an urgent need to evaluate LLMs’ healthcare values before their widespread clinical deployment [10, 19].
Related Work. Recent efforts have introduced benchmarks for evaluating LLMs’ values, e.g., ValueBench [32], WorldValuesBench [48], Flames [22]. However, they primarily focus on general moral reasoning, typically grounded in the “Helpful, Honest, Harmless” (3H) framework, rather than healthcare ethics. In contrast, the existing medical benchmarks such as MedQA [24], MedExQA [25], and MedBench [10] target professional knowledge evaluation through diagnostic reasoning and medical Q&A. However, none of them evaluate whether LLMs can understand or align with ethical values in nursing contexts.
@@
To fill this gap, we introduce Real-World Nursing Values (NurValues), the first benchmark for nursing value alignment, grounded in five core dimensions derived from international nursing codes issued by the organizations such as the American Nurses Association (ANA) [5], the Nursing and Midwifery Council (NMC) [27], the Nursing Council of Hong Kong (NCHK) [28], and the Chinese Nursing Association (CNA) [12]. The five value dimensions - Altruism, Human Dignity, Integrity, Justice, and Professionalism - represent widely accepted ethical principles in nursing practice. We construct the benchmark through a five-month longitudinal field study conducted across three hospitals at different levels (primary, secondary, and tertiary). By shadowing nurses in their daily routines and systematically documenting their behaviors, we collect 1,100 real-world nursing behavior instances.
@@
Then, each of the 1,100 instances is annotated by five licensed clinical nurses, who annotate the nursing value(s) expressed in the instance and judge whether the nurse’s behavior aligns with them. To construct a balanced dataset, we then use o1 to generate a counterfactual version for each instance, preserving the original context while reversing the ethical polarity (e.g., “administering two different vaccines to a child in two arms” vs. “injecting both into the same arm”). This yields both value-aligned and value-violating variants for each instance, resulting in a total of 2,200 labeled samples. We refer to this collection as the Easy-Level dataset.
To increase adversarial complexity, we leverage current jailbreaking techniques to transform each instance into a dialogue-based format. While preserving the original behavioral content and value orientation, we embed richer contextual cues and misleading signals such as persuasion, traps, and deception. This results in the Hard-Level dataset, where each dialogue presents the same nursing behavior in a more ambiguous and adversarial context. The final NurValues benchmark includes 2,200 Easy-Level and 2,200 Hard-Level instances. Tab. 1 compares NurValues with other benchmarks.
Table 1: Comparison of NurValues with other benchmarks.
Dataset Size Language Values? Hierarchical? Medical? Real-World?
MedQA [24] 61.1k English, Chinese $\times$ $\times$ $\checkmark$ $\checkmark$
MedBench [10] 300k Chinese $\times$ $\times$ $\checkmark$ $\checkmark$
MedSafetyBench [19] 1.8k English $\checkmark$ $\times$ $\checkmark$ $\times$
ValueBench [32] 2.0k English $\checkmark$ $\times$ $\times$ $\times$
Flames [22] 2.2k Chinese $\checkmark$ $\times$ $\times$ $\times$
ETHICS [20] 130k English $\checkmark$ $\times$ $\times$ $\checkmark$
Safety-prompts [42] 100k Chinese $\checkmark$ $\times$ $\times$ $\times$
NurValues (Ours) 4.4k English, Chinese $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
We present comprehensive empirical evaluations of NurValues over 23 SoTA LLMs (18 general LLMs and 5 medical LLMs), and compare their results across different dimensions. The results yield three key findings: (1) DeepSeek-V3 achieves the highest performance on the Easy-Level dataset (94.55 in Ma-F1), while Claude 3.5 Sonnet leads on the Hard-Level dataset (89.43 in Ma-F1). Additionally, general LLMs consistently outperform medical LLMs; (2) among the five nursing value dimensions, Justice is consistently the most difficult for LLMs to assess accurately; and (3) in-context learning (ICL) strategies significantly improve model performance. For example, DeepSeek-V3 improves its Ma-F1 from 34.29 to 57.32 with CoT prompting on Hard-Level dataset. This demonstrates the effectiveness and potential of the proposed benchmark. Our main contributions are written as:
@@
  • We propose NurValues, the first real-world nursing value benchmark, based on a five-month field study across three hospitals. We hope that NurValues can advance the value alignment of LLMs in healthcare settings.
  • It covers five core value dimensions, which are Altruism, Human Dignity, Integrity, Justice, Professionalism (see Fig. 4 in App. A), and includes two difficulty levels: an Easy-Level and a Hard-Level subset.
  • We conduct systematic experiments on 23 LLMs to evaluate their strengths and limitations.
@@ 2 The NurValues Benchmark
Definition of Nursing Values. There is currently no universally accepted definition of nursing values. However, according to guiding documents from internationally recognized organizations such as the International Council of Nurses (ICN) [23], nursing values typically refer to the core ethical principles and professional beliefs that nurses are expected to uphold in clinical practice. Despite differences in healthcare systems, cultural norms, and regulatory frameworks across countries, the essence of nursing remains consistent: to provide patient-centered care that is safe, respectful, equitable, and professional.
To construct a value framework that both captures international consensus and provides practical guidance, we conduct a systematic review of several authoritative ethical codes, including those issued by the ANA, the NMC, the NCHK, the CNA, and the ICN, supplemented by relevant literature [36, 34, 21]. Then, we identify a set of core value dimensions that are commonly recognized across institutions, excluding those not universally endorsed. Finally, we derive five core value dimensions that form the ethical backbone of our benchmark:
@@
  • Justice involves fair and equitable treatment for all patients, regardless of their background or circumstances. It includes nondiscrimination, equal access to care, and the ethical distribution of medical resources.
  • Integrity encompasses honesty, consistency, ethical decision-making, and accountability. It requires nurses to adhere to professional standards, follow institutional protocols, and take responsibility for their actions.
@@ 2.1 Data Acquisition
(2) Coverage of 11 clinical departments. To ensure scenario diversity, we gather data from 11 frontline clinical departments, organized into four categories: (1) Emergency and Critical Care Units: Emergency Room, Intensive Care Unit (ICU), and Department of Respiratory and Critical Care Medicine. (2) Chronic Care Units: Hematology, Geriatrics, and Rehabilitation Medicine. (3) General Care Units: General Medicine, Orthopedics, and Obstetrics. (4) Specialized Services: Infusion Hall and Vaccination Center.
This structure ensures the dataset encompasses a wide range of nursing tasks: from routine care to acute emergency response, from preventive services to long-term chronic care—thereby enriching its behavioral and ethical diversity.
(3) Five-month longitudinal observation. From December 2024 to April 2025, we conduct a five-month longitudinal field study (proofs please see Fig. 5). We enter the hospitals 2-3 times per week, for approximately 6 hours per session, accumulating over 300 hours of on-site observation. Using a non-intrusive, observer-as-witness methodology, we follow nurses across outpatient clinics, inpatient wards, and emergency units, systematically recording verbal interactions, clinical decision-making, and communicative behaviors involving patients, families, physicians, and colleagues.
@@
(5) Privacy protection. This study involves the collection of real-world behavioral data from clinical nursing settings. All data were collected through non-intrusive field observation in three hospitals. No personally identifiable information (PII) was recorded, and all interactions were anonymized during transcription. Participants, including nurses and other medical personnel, were informed of the purpose, scope, and non-intrusive nature of the study. Verbal informed consent was obtained when necessary, following the ethical review procedures of the affiliated institution.
@@ 2.2 Dataset Construction
Figure 1: The pipeline for dataset construction.
@@
Step 1: Expert annotation. Based on the 976 instances collected from hospitals, we invite five clinically experienced experts (including 1 head nurse and 4 registered nurses) to participate in the annotation task. For each instance, four annotators independently select up to two of the five predefined nursing value dimensions, and judge whether the behavior aligns with the selected value(s). The fifth expert serves as a reviewer: instances with unanimous agreement among the four annotators are accepted directly. For the case with disagreements, she re-annotates the instance and determines the golden label. To guarantee high-quality annotations, we calculate the Fleiss’ kappa score, $\kappa=0.73$, which means the five experts have reached high agreement. further details on human annotation procedure, please see App. C. Among all annotated instances, 124 are labeled with two distinct value dimensions. We duplicate each of these samples, associating the same text with two different value labels respectively. The remaining 852 instances are assigned a single value label. This yields a total of 1,100 annotated samples.
Step 2: Counterfactual augmentation. To construct a balanced dataset, we employ value-flipping augmentation using o1. For each annotated instance, we generate a counterfactual instance by reversing its value polarity, by turning a value-aligned behavior into a value-violating one, or vice versa, while keeping the original context, length, and factual integrity unchanged. For example, “administering two different vaccines to a child in two separate arms” is flipped into “injecting both vaccines into the same arm”. All counterfactuals are manually reviewed and revised as necessary. This process results in the Easy-Level dataset containing 2,200 samples, as shown in Tab. 2).
@@
Step 3: Difficulty escalation. To increase adversarial complexity, we leverage current jailbreaking techniques for LLMs and o1 to rewrite each instance as a dialogue between two virtual speakers. Each dialogue begins with an accurate restatement of the original case, followed by a persuasive exchange in which one speaker attempts to manipulate the other’s ethical judgment through reasoning traps, biased framing, or plausible but misleading justification. While retaining the original behavioral scenario and value label, the dialogue format introduces subtle misleading signals and contextual noise to obscure moral clarity. All generated dialogues are manually verified to ensure semantic fidelity and consistency with the original scenario. Hence, the Hard-Level subset also has 2,200 samples. This dual-layer benchmark enables systematic evaluation of LLMs’ alignment with nursing values under both standard and adversarial conditions.
@@ 2.3 Dataset Analysis
Table 2: Value dimensions distribution in Easy/Hard-Level datasets. Parentheses show abbreviations.
Dimension Positive Negative Sum
Altruism (ALT) 239 239 478
Human Dignity (HD) 266 266 532
Integrity (INT) 261 261 522
Justice (JUS) 37 37 74
Professionalism (PRO) 297 297 594
Total 1100 1100 2200
The distribution of value dimensions. Table 2 reports the distribution of samples across the five nursing value dimensions in both the Easy- and Hard-Level subsets. Professionalism is the most represented value, accounting for 594 instances (27.0%), followed by Human Dignity (532), Integrity (522), and Altruism (478). In contrast, Justice is notably underrepresented, with only 74 samples in total (3.36%). This imbalance reflects the relative rarity of justice-related behaviors in the observed real-world nursing scenarios.
Distribution of nursing behavior types. We manually categorize the 1,100 real-world cases based on the nature of the nurse’s action (Fig 2 A). The cases fall into five major categories: (1) Operational: technical procedures such as injections, transfusions, wound care, or catheter placement (38.1%); (2) Communicative: verbal interactions, health education, counselling, or emotional reassurance (36.7%); (3) Emergency: time-critical or high-risk interventions, rapid assessment, or resuscitation measures (7.5%); (4) Support & Management: resource preparation, shift hand-over, family coordination, or logistical support (5.3%); (5) Other: cases that do not clearly fit the above categories or contain multiple mixed actions (12.4%).
Operational (38.1%) and Communicative (36.7%) cases together account for roughly two-thirds of the corpus, underscoring that routine nursing practice is dominated by technical procedures and nurse–patient communication. Although Emergency scenarios represent only 7.5% of the data, they concentrate the highest ethical stakes and thus serve as a stress-test for model safety. The Support & Management category (5.3%) covering shift hand overs, supply preparation, and family coordination, captures indirect yet indispensable activities. The remaining 12.5% of mixed or ambiguous cases highlight the complexity of real-world nursing contexts.
Topic consistency across Easy- and Hard-Level datasets. As shown in Fig. 2 B, our evaluation confirms that the rewritten dialogues retain high topical consistency with the original instances. Using three SoTA LLMs to score semantic similarity, over 82% of instances receive a score $\geq$ 8, and all LLMs yield average scores above 8, showing that the Hard-Level subset increases reasoning complexity without harming the original semantic structure or ethical context (for details, see App. D).
Figure 2: A:The topic consistency between the simple instances and the complicated dialogues across three LLMs. B: The distribution of nursing behavior types.
Comparison across Easy- and Hard-Level datasets. We perform a statistical analysis of linguistic features across the Easy- and Hard-Level datasets. Specifically, we examine four key metrics: mean
Table 3: The linguistic features analysis.
Lang. Dataset M Len. M TTR G TTR M Clauses Count
EN Easy- 24.07 0.83 0.08 0.89
Hard- 235.79 0.63 0.03 8.42
CN Easy- 32.01 0.91 0.07 0.54
Hard- 440.28 0.55 0.02 4.26
length of text (M Len.), mean type-token ratio (M TTR), global type-token ratio (G TTR), and mean clauses count (M Clauses Count). As shown in Tab. 3, both the Chinese and English versions demonstrate that Easy-Level samples are generally shorter, use simpler sentence structures, and exhibit lower lexical diversity than their Hard-Level counterparts. These differences confirm that the Hard-Level subset presents greater linguistic complexity and cognitive load, consistent with its design purpose.

### v2

@@ Abstract
While LLMs have demonstrated medical knowledge and conversational ability, their deployment in clinical practice raises new risks: patients may place greater trust in LLM-generated responses than in nurses’ professional judgments, potentially intensifying nurse–patient conflicts. Such risks highlight the urgent need of evaluating whether LLMs align with the core nursing values upheld by human nurses. This work introduces the first benchmark for nursing value alignment, consisting of five core value dimensions distilled from international nursing codes: Altruism, Human Dignity, Integrity, Justice, and Professionalism. We define two-level tasks on the benchmark, considering the two characteristics of emerging nurse–patient conflicts. The Easy-Level dataset consists of 2,200 value-aligned and value-violating instances, which are collected through a five-month longitudinal field study across three hospitals of varying tiers; The Hard-Level dataset is comprised of 2,200 dialogue-based instances that embed contextual cues and subtle misleading signals, which increase adversarial complexity and better reflect the subjectivity and bias of narrators in the context of emerging nurse-patient conflicts. We evaluate a total of 23 SoTA LLMs on their ability to align with nursing values, and find that general LLMs outperform medical ones, and Justice is the hardest value dimension. As the first real-world benchmark for healthcare value alignment, NurValues provides novel insights into how LLMs navigate ethical challenges in clinician–patient interactions.  [^1] The dataset and the code are available at https://github.com/BenYyyyyy/NurValues
Figure 1: Adverse consequences of misalignment between LLM and human nursing values. Left: when a patient loses independence, a nurse must prioritize the patient’s safety and access to treatment, but the LLM misjudged this as a failure to respect the patient’s dignity. Right: when a patient autonomously refused, the nurse should respect the decision. However, the LLM misinterpreted the nurse’s disregard of the patient’s will as altruism, leading to adverse consequences for the patient.
@@ 1 Introduction
Through post-training on synthetic medical data, large language models (LLMs) have gained the understanding and conversational ability in the medical domain [46, 31, 44]. In clinical practices, however, these models still face challenges in real clinical settings, among which tensions between healthcare professionals and patients are particularly pronounced [33]. Owing to their wide range of duties and frequent interactions with patients, nurses have long been at the forefront of such conflicts [50, 5]. Unlike the knowledge-based disagreements that often arise in physicians’ diagnostic practices, nursing conflicts are more frequently rooted in the differences of value judgments.
With the emergence of LLM-assisted nursing practices, the conflicts have appeared in new forms: patients may tend to place greater trust in responses generated by LLMs [43]. When such responses diverge from nurses’ professional judgments, tensions may be further amplified [23]. This shift in trust not only risks undermining the professional authority of nurses but also intensifies nurse–patient conflicts in subtle ways. Therefore, it is crucial to explore whether LLMs can embody values consistent with nursing practice. Insufficient value alignment may turn LLMs into new sources of conflict, whilst well-aligned LLM responses can alleviate conflicts and foster trust. In this context, there is an urgent need to establish evaluation benchmarks for evaluating the alignment of LLMs with the nursing values upheld by human nurses.
Fig. 1 presents two cases that illustrate the potential risks of value misalignment between LLMs and nurses in nursing contexts. These cases highlight distinctive characteristics absent in traditional conflicts, which extend beyond the scope of existing medical or value assessment benchmarks.
  • First, existing value benchmarks (e.g., ValueBench [39], WorldValuesBench [55], Flames [25]) focus on general moral reasoning under the “Helpful, Honest, Harmless” (3H) framework, while medical benchmarks (e.g., MedQA [29], MedExQA [30], MedBench [11]) mainly test diagnostic knowledge. Neither evaluates whether LLMs align with ethical principles in nursing, leaving a critical gap.
  • Second, in many nursing scenarios, conflicts are not triggered by patients’ direct and explicit requests, but rather by accounts relayed through patients or their family members. Such accounts are frequently infused with positional biases, emotional embellishments, or intentional and unintentional framing. LLMs must extract factual information from ambiguous narratives and render judgments grounded in nursing values. This is fundamentally different from existing medical safety evaluations (e.g., MedSafetyBench [22]), which primarily rely on straightforward “harmful request—response” matching. Hence, evaluating nursing value alignment requires the capacity for complex ethical reasoning in uncertain and subjective contexts.
@@
To fill the gaps, we introduce Real-World Nursing Values (NurValues), the first benchmark for nursing value alignment, grounded in five core dimensions derived from international nursing codes issued by the authoritative international organizations, including the American Nurses Association (ANA) [6], the Nursing and Midwifery Council (NMC) [34], the Nursing Council of Hong Kong (NCHK) [35], and the Chinese Nursing Association (CNA) [13]. The five value dimensions - Altruism, Human Dignity, Integrity, Justice, and Professionalism - represent widely accepted ethical principles in nursing practice. To support fair and comprehensive LLM evaluation on key nursing value dimensions, we meticulously designed a two-level benchmark, which can address the two characteristics of new nurse–patient conflicts outlined above.
@@
  • (1) We conducted a five-month field study in three hospitals (primary, secondary, and tertiary), systematically collecting 1,100 real-world cases from nurses’ daily work, which involve both clinical knowledge and nursing values. Then, each of the 1,100 instances is annotated by five licensed clinical nurses, who annotate the nursing value(s) expressed in the instance and judge whether the nurse’s behavior aligns with them. To construct a balanced dataset, we then use OpenAI o1 to generate a counterfactual version for each instance, preserving the original context while reversing the ethical polarity (e.g., “administering two different vaccines to a child in two arms” vs. “injecting both into the same arm”). The generated counterfactual samples, verified by domain experts and combined with the original instances, constitute a total of 2,200 labeled samples, which together form the Easy-Level dataset.
  • (2) Real high-conflict and emotionally charged dialogues often contain sensitive expressions, extensive personal information and chaotic interactions, making it difficult to collect and release these dialogues in the raw form. Therefore, to better reflect such conflicts under LLM involvement, we transform each authentic case in the Easy-Level dataset into a dialogue format enriched with contextual cues and misleading signals (e.g., persuasion, traps, deception). This approach preserves realism while ensuring systematic coverage of complex, subjective scenarios, ultimately yielding a more challenging Hard-Level dataset of 2,200 samples.
@@
We present comprehensive empirical evaluations of NurValues over 23 SoTA LLMs. The results yield three key findings: (1) DeepSeek-V3 achieves the highest performance on the Easy-Level dataset (94.55 in Ma-F1), while Claude 3.5 Sonnet leads on the Hard-Level dataset (89.43 in Ma-F1) that showcases its excellent professional nursing knowledge and ethical reasoning skills. Additionally, general LLMs consistently outperform medical LLMs; (2) among the five nursing value dimensions, Justice is consistently the most difficult for LLMs to assess accurately; and (3) in-context learning (ICL) strategies significantly improve model performance. This demonstrates the effectiveness and potential of the proposed benchmark. Our main contributions are:
@@
  • We present NurValues, the first real-world nursing value benchmark, constructed from a five-month field study across three hospitals, to evaluate the alignment of LLMs with professional nursing values.
  • NurValues covers five core nursing value dimensions (Fig. 4 in App. A) and is organized into Easy-Level and Hard-Level, enabling evaluation under both standard and complex ethical contexts.
  • We conducted systematic experiments on 23 LLMs to evaluate their strengths and limitations on tasks measuring alignment with nursing values, providing initial insights for future model improvements and comparisons in terms of nursing values.
@@ 2 The NurValues Benchmark
Definition of Nursing Values. Nursing values, though lacking a universal definition, generally encompass core ethical principles and professional beliefs guiding safe, respectful, equitable, and patient-centered care across diverse contexts [27]. Through a systematic review of major ethical codes (ANA, NMC, NCHK, CNA, and ICN) and relevant literature [42, 40, 24], we identify and distill five universally recognized core value dimensions that form the ethical backbone of our benchmark:
@@
  • Justice involves fair and equitable treatment for patients, regardless of their circumstances. It includes nondiscrimination, equal access to care, and the ethical distribution of medical resources.
  • Integrity encompasses honesty, consistency, ethical decision-making, and accountability. It requires nurses to adhere to standards, follow institutional protocols, and take responsibility for their actions.
@@ 2.1 Data Acquisition
(2) Coverage of 11 clinical departments. To ensure scenario diversity, we gather data from 11 frontline clinical departments, organized into four categories: (1) Emergency and Critical Care Units: Emergency Room, Intensive Care Unit (ICU), and Department of Respiratory and Critical Care Medicine. (2) Chronic Care Units: Hematology, Geriatrics, and Rehabilitation Medicine. (3) General Care Units: General Medicine, Orthopedics, and Obstetrics. (4) Specialized Services: Infusion Hall and Vaccination Center. This structure ensures the dataset encompasses a wide range of nursing tasks: from routine care to acute emergency response, from preventive services to long-term chronic care—thereby enriching its behavioral and ethical diversity.
(3) Five-month longitudinal observation. From December 2024 to April 2025, we conduct a five-month longitudinal field study. We enter the hospitals 2-3 times per week, for approximately 6 hours per session, accumulating over 300 hours of on-site observation. Using a non-intrusive, observer-as-witness methodology, we follow nurses across outpatient clinics, inpatient wards, and emergency units, systematically recording verbal interactions, clinical decision-making, and communicative behaviors involving patients, families, physicians, and colleagues.
@@
(5) Privacy protection. This study involves the collection of real-world behavioral data from clinical nursing settings. All data were collected through non-intrusive field observation in three hospitals. No personally identifiable information (PII) was recorded, and all interactions were anonymized during transcription. Participants, including nurses and other medical personnel, were informed of the purpose, scope, and non-intrusive nature of the study. The study has been reviewed and approved by the hospital ethics committees.
@@ 2.2 Dataset Construction
Figure 2: The pipeline for dataset construction. The Label in each step specifies the nursing values that the case or dialogue is related to. The 0/1 indicator shows whether the nurse’s behavior violates or aligns with the corresponding nursing value.
@@
Step 1: Expert annotation. Based on the 976 instances collected from hospitals, we invite five clinically experienced experts (including 1 head nurse and 4 registered nurses) to participate in the annotation task. All five nurses received adequate training and clear annotation guidelines before participating in the study. For each instance, four annotators independently select up to two of the five predefined nursing value dimensions, and judge whether the behavior aligns with the selected value(s). The fifth expert serves as a reviewer: instances with unanimous agreement among the four annotators are accepted directly. For the case with disagreements, the fifth expert re-annotates the instance and determines the golden label. To guarantee high-quality annotations, we calculate the Fleiss’ kappa score, $\kappa=0.73$, which means the five experts have reached high agreement. Further details on human annotation procedure, please see App. C. Among all annotated instances, 124 are labeled with two distinct value dimensions. We duplicate each of these samples, associating the same text with two different value labels respectively. The remaining 852 instances are assigned a single value label. This yields a total of 1,100 annotated samples.
Step 2: Counterfactual augmentation. To construct a balanced dataset, we employ value-flipping augmentation using o1. For each annotated instance, we generate a counterfactual instance by reversing its value polarity, by turning a value-aligned behavior into a value-violating one, or vice versa, while keeping the original context, length, and factual integrity unchanged. For example, “administering two different vaccines to a child in two separate arms” is flipped into “injecting both vaccines into the same arm”. All counterfactuals are manually reviewed and revised as necessary. This process results in the Easy-Level dataset containing 2,200 samples, as shown in Fig. 3 A.
@@
Step 3: Difficulty escalation. To reflect the subjectivity and bias of the narrator in new doctor-patient conflicts, we leverage jailbreaking techniques for LLMs and o1 to rewrite each instance as a dialogue between two virtual speakers. Each dialogue begins with an accurate restatement of the original case, followed by a persuasive exchange in which one speaker attempts to manipulate the other’s ethical judgment through reasoning traps, biased framing, or plausible but misleading justification. While retaining the original behavioral scenario and value label, the dialogue format introduces subtle misleading signals and contextual noise to obscure moral clarity. All generated dialogues are manually verified to ensure semantic fidelity and consistency with the original scenario. Hence, the Hard-Level subset also has 2,200 samples. This dual-layer benchmark systematically evaluates LLMs’ alignment with nursing values under standard and adversarial conditions.
@@ 2.3 Dataset Analysis
The distribution of value dimensions. Fig. 3 A reports the distribution of samples across the five nursing value dimensions in both the Easy- and Hard-Level subsets. Professionalism is the most represented value, accounting for 594 instances (27.0%), followed by Human Dignity (532), Integrity (522), and Altruism (478). In contrast, Justice is notably underrepresented, with only 74 samples in total (3.36%). This imbalance reflects the relative rarity of justice-related behaviors in the observed real-world nursing scenarios.
Topic consistency across Easy- and Hard-Level datasets. As shown in Fig. 3 B, our evaluation confirms that the rewritten dialogues retain high topical consistency with the original instances. For each of the 3 SoTA LLMs, over 82% of instances receive a semantic similarity score $\geq$ 8. And all LLMs yield average scores above 8, showing that the Hard-Level subset increases reasoning complexity without harming the original semantic structure or ethical context. In addition, we invite two human evaluators to score topical consistency on 200 randomly sampled, label-balanced instances. As shown in Fig. 3 C, both evaluators assigned average scores above 7. Considering their stated avoidance of extreme scores in the post-experiment questionnaire, this suggests that our adversarial dialogues successfully preserve the topic of the original sentence.(for details, see App. D).
Figure 3: A: Value dimensions distribution in label-balanced Easy/Hard-Level datasets. Parentheses show abbreviations. B: Topic consistency of simple instances and complicated dialogues annotated using three LLMs on the NurValues dataset. C: Topic consistency of simple instances and complicated dialogues annotated by two humans and three LLMs on 200 randomly sampled instances.